what are the "magic" rules for the resistor, capacitor, and inductor? what is the current in each?
Answer #1

The reistor ,capacitor and inductor all three are passive component.

Current in Resistor is given by

I_R = \frac{V_R}{R}

Current though Capacitor is given by

I_C = C.\frac{\mathrm{d} V_C}{\mathrm{d} t}

and Current though Capacitor is given by

I_L = \frac{1}{L}.\int V_L.dt